I don't get this, 9y-1=6y +5
9y +1 -1 = +1 6y +5 or am I starting out wrong from the begining.

Answer by jessicawhidden(3)

you are doing ok, but i fear you will confuse yourself.
9y-1=6y+5
get your variables to one side
9y-6y-1=6y-6y+5
get your variable alone
3y-1+1=5+1
3y=6
3y/3=6/3
y=2
check your answer by plugging it into the origional equation
9*2-1=6*2+5
18-1 =12+5
17=17